. Find the square of the following numbers.  ii. 35  
l,,kgyv,gyk,yfkty

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student



ii. (35)2
= (30+5 )2
= (30)2+ (5)2+ 2×30×5 =900+25+300
= 1225




Thanks